Shift report 20000630 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Jun 30 2000 - 23:59:00 EDT

  • Next message: http account: "Shift report 20000701 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Jun 30 23:58:59 2000, Shift: 16:00-24:00
    Supervisor: kh, with: bs,ff
    Log book#2 pages 101 - 101
    
    Shift Summary:
    ------------------------------------------------------------
    No beam during most of the shift until 23:35 when the
    monitor suddenly appeared with 
    "beam cogged at gamma=70"
    
    At that point, we turned on voltages on tpc's and started a
    run.  Beam intensity is 215/292.  Data rate into computer
    is very slow.
    
    Problem with MTPC1 drift.  It always trips.  Can never
    get to -4980.  Tried with lower voltage, but it even trips
    at 4500.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Jul 01 2000 - 00:02:59 EDT